OpenStudy (anonymous):

Find all replacements (less than the modulus. for the question mark that make the statement true. 7 + ? ≡ 3 (mod 8)

OpenStudy (anonymous):

what they mean by less than modulus ? less than 8 maybe ha?

OpenStudy (anonymous):

Lol maybe you don't understand this type of math?

OpenStudy (anonymous):

maybe...

OpenStudy (anonymous):

but i wanna learn it...what they mean by that...

OpenStudy (anonymous):

Well im not to good in this, that's why im asking the question.

OpenStudy (amistre64):

another way to write 3 mod8 is: 8n + 3

OpenStudy (anonymous):

n is 7 im supposing?

OpenStudy (amistre64):

n is usually just a placeholder of an integer value in this case 7 + x = 8n + 3 x = 8n + 3 - 7 x = 8n - 4

OpenStudy (amistre64):

if you run thru your 8s times tables, and add 3 to each one; you get the set of 3 mod8

OpenStudy (amistre64):

-4 -3 -2 -1 0 1 2 3 4 5 6 7 8 it looks like this setup is the same as saying x = 8x + 5 ; or x = 5 mod8

OpenStudy (amistre64):

i almost typed that out without a mistake :)

OpenStudy (anonymous):

Nice, thanks.

OpenStudy (amistre64):

youre welcome but as the first poster said, it not clear with the "less than the modulus" actually means. If i were to take a gander; it would prolly define the setup x = 8n + 5 such that n < 1